# Solvable Tuple Patterns and Their Applications to Program Verification

### v1

@@ 3.2 Direct Applications of STP Inference
[…]
If $T\cup\{t_{\mathit{new}}\}$ satisfies $D$, then we add $t_{\mathit{new}}$ to $T$, and check whether $T$ is strong enough to imply the goal $G$ (lines 9-10). If so, we can conclude that the CHCs $S$ is satisfiable, and return $T$ as a witness. If $T\models G$ does not hold, then we go back to the beginning of the loop to find another STP.
If $t_{\mathit{new}}$ belongs to $T$ or $T\cup\{t_{\mathit{new}}\}\models D$ does not hold, then we collect more samples (line 13) and go back to the beginning of the loop. There are at least two reasonable methods to collect more samples. One is to apply the same procedure as $\mathtt{collect\_samples}$. The other method, which is applicable when $T\cup\{t_{\mathit{new}}\}\models D$ does not hold, is to generate new samples in a counterexample manner: pick a clause $d\in D$ such that $T\cup\{t_{\mathit{new}}\}\models d$ does not hold, and obtain a counterexample by using an SMT solver. In the latter method, there is actually no guarantee that the found value $\widetilde{v}$ belongs to the least model $m_{D,P}$; thus it must be filtered out on line 5 (see Remark 3.2).
Remark 3.2.
If the procedure $\mathtt{collect\_more\_samples}$ generates new samples in the counterexample-guided manner (from an unsatisfied clause $d$) as mentioned above, the generated samples may not belong to the least model of $D$. For example, recall the running example in Example 3.1. Suppose $T=\emptyset$ and $t_{\mathit{new}}=(l_{1},l_{2},l_{1}l_{2})$. Then, $T\cup\{t_{\mathit{new}}\}$ does not satisfy the clause (3), i.e., the following formula is invalid:
$\forall x,l_{1}^{\prime},l_{2},l_{3}.l_{1}^{\prime}xl_{2}=l_{3}\Rightarrow xl_{1}^{\prime}l_{2}=l_{3}.$
Suppose that an SMT solver generates $x=0,l_{1}^{\prime}=1\cdot 2,l_{2}=\epsilon,l_{3}=1\cdot 2\cdot 0$ as a counterexample. Then, $(x\cdot l_{1}^{\prime},l_{2},l_{3})=(0\cdot 1\cdot 2,\epsilon,1\cdot 2\cdot 0)$ would be generated as a new sample, but it does not belong to the least model $\{(l_{1},l_{2},l_{3})\mid l_{3}=l_{1}^{R}l_{2}\}$. Adding this sample would result in the failure to find the appropriate STP $(l_{1},l_{2},l_{1}^{R}l_{2})$. ∎
We now discuss properties of the algorithm. The soundness of the procedure is obvious.
Theorem 3.2 (soundness).
Let $D$ be a set of definite clauses, $G$ be a set of goal clauses, and $S=D\cup G$. Then we have:
  1. 1.
    If $\textsc{solve}(D,G)$ returns SAT($T$), then $S$ is satisfiable.
  2. 2.
    If $\textsc{solve}(D,G)$ returns UNSAT, then $S$ is unsatisfiable.
Proof.
The first property is trivial as the procedure returns SAT($T$) only when $T\models D$ and $T\models G$. To see the second property, notice that the predicate variable occurs only positively in $\neg G$. Thus, $\neg G$ is monotonic with respect to the interpretation of $P$. Since $\mathtt{true\_samples}(M)$ is a subset of the least model of $D$, $\mathtt{true\_samples}(M)\models\neg G$ implies that $\neg G$ is also satisfied by the least model. By Fact 3.1, $S$ is unsatisfiable. ∎
The completeness (the property that if $S$ is satisfiable, then $\textsc{solve}(D,G)$ eventually returns SAT($T$)) does not hold for the obvious reason that there exists a system of CHCs whose model cannot be expressed by using STPs. Relative completeness, however, holds under certain assumptions, as described below.
Theorem 3.3.
Let $D$ be a set of definite clauses, $G$ be a set of goal clauses, and $S=D\cup G$. Let $\{P\mapsto m_{D,P}\}$ be the least model of $D$. Assume the following properties on subprocedures of solve.
  • $\mathtt{STPinf}(M)$ returns an STP for $M$ in a fair manner; i.e., if $\mathtt{STPinf}$ is called infinitely often, each time with a possibly different $M$ such that $(\widetilde{x},[\widetilde{x}/M])\longrightarrow^{*}(t,\Theta)$, then $t$ is eventually returned.
  • $\mathtt{collect\_more\_samples}{}(D,t_{\mathit{new}})$ on line 13 returns only a subset of $m_{D,P}$ and it is also fair, in the sense that if $\widetilde{v}$ belongs to the least model of $D$ and $\mathtt{collect\_more\_samples}$ is called infinitely often, then it eventually returns a set containing $\widetilde{v}$.
Suppose that there exists an STP $t$ such that $m_{D,P}\models_{s}t$ and $\{P\mapsto\mathcal{L}(t)\}$ is a model of $S$. Then, $\textsc{solve}(D,G)$ eventually outputs $\textrm{SAT}(T)$ for some $T$.
In the statement above, $m_{D,P}\models_{s}t$ is an extension of $M\models_{s}t$ for a possibly infinite set: for $m=\{\widetilde{v}_{i}\mid i\in\omega\}$ (where it may be the case $\widetilde{v}_{i}=\widetilde{v}_{j}$ for some $i\neq j$), $m\models_{s}t$ means that there exists a family of substitutions $\theta_{i}(i\in\omega)$ such that (i) $\widetilde{v}_{i}=\theta_{i}t$, and (ii) there exists no variable $x$ such that $\theta_{i}(x)=\epsilon$ for all $i\in\omega$. We have the following lemma, which is a strengthened form of Theorem 2.4.
Lemma 3.4.
Let $\widetilde{x}=x_{1},\ldots,x_{n}$ be mutually distinct variables, and $m\subseteq(\Sigma^{*})^{n}$. If $m\models_{s}t$ and $t\in\mathcal{T}_{n}$, then there exists a finite subset $M^{\prime}$ of $m$ such that, for all $M$ that satisfies $M^{\prime}\subseteq M\subseteq m$, $(\widetilde{x},[M/\widetilde{x}])\longrightarrow^{*}(t,\Theta)$ holds for some $\Theta$. [^9] Recall that tuple patterns are identified up to variable renaming.
Proof.
The rewriting relations can be naturally extended to the case where $M$ is an infinite set (in other words, the numbers of rows of $M$ is infinite). Thus, by Theorem 2.4, we have $(\widetilde{x},[m/\widetilde{x}])\longrightarrow^{*}(t,\Theta)$ for some $\Theta$. In each application of R-Prefix, etc. where the side condition $M[*][i]\neq\emptyset$ exists, pick a row index $j$ such that $M[j][i]\neq\emptyset$. Let $I$ be the set of such row indices (which is finite), and let $M^{\prime}$ be the subset of $m$ consisting of the rows indexed by $I$. Then, for any $M$ such that $M^{\prime}\subseteq M\subseteq m$, the same reduction sequence is applicable to $M$; hence we have $(\widetilde{x},[M/\widetilde{x}])\longrightarrow^{*}(t,\Theta)$ as required. ∎
We are now ready to prove Theorem 3.3.
Proof of Theorem 3.3.
Suppose $m_{D,P}\models_{s}t$ and $\{P\mapsto\mathcal{L}(t)\}$ is a model of $S$. By the soundness, $\textsc{solve}(D,G)$ cannot return UNSAT. Thus, $\textsc{solve}(D,G)$ either returns $\textrm{SAT}(T)$ for some $T$, or falls into an infinite loop. We show the latter cannot happen. For each $M$, the set $\{t^{\prime}\in\mathcal{T}\mid M\models_{s}t^{\prime}\}$ is finite. Therefore, if the body of the while-loop is executed infinitely often, $\mathtt{collect\_more\_samples}$ on line 13 is called infinitely often. By the assumption on $\mathtt{collect\_more\_samples}$, every element of $m_{D,P}$ is eventually added to $M$. By the assumption $m_{D,P}\models_{s}t$ and Lemma 3.4, there exists a finite subset $M_{0}$ of $m_{D,P}$ such that $M_{1}\models_{s}t$ holds for any $M_{0}\subseteq M_{1}\subseteq m_{D,P}$. Since $M$ eventually contains all the elements of $M_{0}$, we have $M\models_{s}t$. At this point, by Theorem 2.4, $\mathtt{STPinf}{}$ eventually outputs $t$ and terminates. ∎
Remark 3.3.
In the assumption on $\mathtt{STPinf}$, we do not require $(t,\Theta)\;\mathbin{\ \not\!\!\longrightarrow}$. If we weaken the assumption on $\mathtt{STPinf}$ by adding such a condition, then the relative completeness would not hold. In fact, consider the following set $S$ of CHCs:
$P(0\cdot 0,0\cdot 0).\qquad P(1\cdot 1,1\cdot 1).\qquad P(x\cdot l_{1},l_{2})\land x>1\Rightarrow P(l_{1},l_{2}).\qquad P(l_{1},l_{2})\land\mathit{false}\Rightarrow\mathit{false}.$
$S$ is obviously satisfiable, with the model $P(l_{1},l_{2})\equiv\mathit{true}$ (which is representable by the trivial pattern $(x,y)$). However, the least model of the definite clauses is $\{P\mapsto M\}$ with $M=\{(0\cdot 0,0\cdot 0),(1\cdot 1,1\cdot 1)\}$ and $t$ such that $((l_{1},l_{2}),[(l_{1},l_{2})/M])\longrightarrow^{*}(t,\Theta)\,\;\mathbin{\ \not\!\!\longrightarrow}$ is only $(l_{1},l_{1})$ (up to renaming). But $\{P\mapsto\mathcal{L}(l_{1},l_{1})\}$ does not satisfy the third clause. ∎
@@ 3.3 Combination with CHC Solving for Integer Arithmetic
As mentioned in Section 1, the tuple patterns alone are often not sufficient to express appropriate invariants. Let us recall the second example (involving take and drop) in Section 1. It can be expressed as the satisfiability of the following CHCs:
$\mathit{Take}(0,l,[]).\qquad n\neq 0\Rightarrow\mathit{Take}(n,[],[]).\qquad n\neq 0\land\mathit{Take}(n-1,l^{\prime},r)\Rightarrow\mathit{Take}(n,x::l^{\prime},x::r).$
$\mathit{Drop}(0,l,l).\qquad n\neq 0\land\mathit{Drop}(n,[],[]).\qquad n\neq 0\land\mathit{Drop}(n-1,l^{\prime},r)\Rightarrow\mathit{Drop}(n,x::l^{\prime},r).$
@@
Here, we have omitted the clauses for $\mathit{Append}$. By the STP inference, we obtain the candidate models: $\mathit{Take}(n,l_{1},l_{2})\equiv\exists l_{3}.l_{1}=l_{2}l_{3}$, $\mathit{Drop}(n,l_{1},l_{2})\equiv\exists l_{3}.l_{1}=l_{3}l_{2}$, and $\mathit{Append}(l_{1},l_{2},l_{3})\equiv l_{3}=l_{1}l_{2}$, but they are not strong enough to derive the goal clause (the last clause above).
To this end, we use a CHC solver over integer arithmetic to refine the candidate models. We first abstract lists to their lengths and obtain the following abstract version of the CHCs.
@@
In general, given a system of CHCs over lists $S=D\cup G$, (i) the list constructors $[]$ and $\mathbin{::}$ are replaced with their length abstractions 0 and $\lambda(x,y)=1+y$, (ii) the list equalities (which implicitly occur in the clause bodies) are replaced with integer equalities; and (iii) the list inequalities $l\neq l^{\prime}$ in the body of goal clauses $G$ are replaced with the integer inequalities, and those in the bodies of definite clauses $D$ are replaced with $\mathit{true}$.
The above abstraction ensures that a model of the abstract version $D^{\prime}$ of definite clauses yields that of the original definite clauses $D$: if $P^{\prime}(\widetilde{x},l^{\prime}_{1},\ldots,l^{\prime}_{m})\equiv\varphi$ satisfies $D^{\prime}$, then $P(\widetilde{x},l_{1},\ldots,l_{m})\equiv[\mathit{len}(l_{1})/l^{\prime}_{1},\ldots,\mathit{len}(l_{m})/l^{\prime}_{m}]\varphi$ satisfies $D$ (where we assume $\widetilde{x}$ are integer arguments and $l_{1},\ldots,l_{m}$ are list arguments); this follows from the soundness of the catamorphism-based abstraction of CHCs [30].
In contrast, there is no such guarantee for the goal clauses: because of the replacement of $_{\mathit{list}}$ and $\neq_{\mathit{list}}$ with $_{\mathit{int}}$ and $\neq_{\mathit{int}}$, the abstract version of goal clauses is neither stronger nor weaker than the original goal clauses. (Notice that $\mathit{len}(l_{1})=\mathit{len}(l_{2})$ is an over-approximation of $l_{1}=l_{2}$, while $\mathit{len}(l_{1})\neq\mathit{len}(l_{2})$ is an under-approximation of $l_{1}\neq l_{2}$ is an underapproximation.) The abstract goal clauses, however, tend to help a CHC solver to find an appropriate invariant. In the case above, if $l\neq l^{\prime}$ were replaced with $\mathit{true}$, we would obtain $\mathit{Take^{\prime}}(n,l,r_{1})\land\mathit{Drop^{\prime}}(n,l,r_{2})\land\mathit{Append^{\prime}}(r_{1},r_{2},l^{\prime})\Rightarrow\mathit{false}$, which is so strong that the abstract CHCs would become unsatisfiable.
@@
$\mathit{Take}(n,l,r)\equiv(\mathit{len}(l)<n\land\mathit{len}(r)=\mathit{len}(l))\lor(\mathit{len}(l)\geq n\land\mathit{len}(r)=n)$
$\mathit{Drop}(n,l,r)\equiv(\mathit{len}(l)<n\land\mathit{len}(r)=0)\lor(\mathit{len}(l)\geq n\land\mathit{len}(r)=\mathit{len}(l)-n).$
@@
$\mathit{Take}(n,l,r)\equiv\exists s.l=rs\land(\mathit{len}(l)<n\land\mathit{len}(r)=\mathit{len}(l))\lor(\mathit{len}(l)\geq n\land\mathit{len}(r)=n)$
$\mathit{Drop}(n,l,r)\equiv\exists s.l=sr\land(\mathit{len}(l)<n\land\mathit{len}(r)=0)\lor(\mathit{len}(l)\geq n\land\mathit{len}(r)=\mathit{len}(l)-n).$
@@
Instead of just using the information obtained from the length abstraction on line 12, it is possible to propagate information gathered from STP inference and length abstraction to each other, as in Nelson-Oppen procedure [41] for theory combinations. For example, suppose that $P(l_{1},l_{2},l_{3})$ is found to imply $l_{3}=l_{1}l_{2}$ by STP inference. Then, each occurrence of $P(l_{1},l_{2},l_{3})$ in the body of a clause can be replaced by $P(l_{1},l_{2},l_{3})\land l_{3}=l_{1}l_{2}$, before applying the length abstraction. Then, a clause of the form $P(l_{1},l_{2},l_{3})\land\cdots\Rightarrow P(l_{4},l_{5},l_{6})$ is abstracted to $P^{\prime}(l_{1},l_{2},l_{3})\land l_{3}=l_{1}+l_{2}\land\cdots\Rightarrow P^{\prime}(l_{4},l_{5},l_{6})$. ∎
@@ 3.4 Beyond Sequence and Integer Constraints
The set/multiset tuple patterns discussed in Section 2.3.4 are also applicable to CHC solving. We informally discuss an application through an example.
@@
Let us consider a system $S_{\mathit{sort}}$ of CHCs consisting of the following clauses:
$\mathit{Insert}(x,[],[x]).\qquad x\leq y\Rightarrow\mathit{Insert}(x,y::l_{2},x::y::l_{2}).$
$x>y\land\mathit{Insert}(x,l_{2},l_{3})\Rightarrow\mathit{Insert}(x,y::l_{2},y::l_{3}).$
$\mathit{Sort}([],[]).\qquad\mathit{Sort}(l_{1},l_{2})\land\mathit{Insert}(x,l_{2},l_{3})\Rightarrow\mathit{Sort}(x::l_{1},l_{3}).$
$\mathit{Sort}(l_{1},l_{2})\land\mathit{Count}(x,l_{1},z_{1})\land\mathit{Count}(x,l_{2},z_{2})\Rightarrow z_{1}=z_{2}.$
Here, $\mathit{Count}(x,l,z)$ means that the number of occurrences of $x$ in $l$ is $z$. So, the last clause means that the insertion sort preserves the multiset of elements.
As discussed in Section 2.3.3, our STP inference is unable to infer that $\mathit{Sort}(l_{1},l_{2})$ implies $l_{2}$ is a sorted list of $l_{1}$. We can, however, prove the satisfiability of the above CHCs by using the extension in Section 2.3.4, as follows. By abstracting lists in the definite clauses to multisets, we obtain:
$\mathit{Insert^{\prime}}(x,\emptyset,{x}).\qquad x\leq y\Rightarrow\mathit{Insert^{\prime}}(x,{y}\uplus l_{2},x\uplus y\uplus l_{2}).$
$x>y\land\mathit{Insert^{\prime}}(x,l_{2},l_{3})\Rightarrow\mathit{Insert^{\prime}}(x,{y}\uplus l_{2},{y}\uplus l_{3}).$
$\mathit{Sort^{\prime}}(\emptyset,\emptyset).\qquad\mathit{Sort^{\prime}}(l_{1},l_{2})\land\mathit{Insert^{\prime}}(x,l_{2},l_{3})\Rightarrow\mathit{Sort^{\prime}}({x}\uplus l_{1},l_{3}).$
Here, by abuse of notation, we identify a singleton multiset with its element, and just write $x\uplus l$ for $\{x\}\uplus l$. By using the STP inference for multisets, we obtain the tuple patterns $(x,y,xy)$ and $(x,x)$ for the arguments of $\mathit{Insert^{\prime}}$ and $\mathit{Sort^{\prime}}$ respectively. Thus, we have a candidate model $\mathit{Sort}(x,y)\equiv\mathit{multiset\_of}(x)=\mathit{multiset\_of}(y)$. Together with $\mathit{Count}(x,l,z)\equiv z=l(x)$, it satisfies the goal clause $\mathit{Sort}(l_{1},l_{2})\land\mathit{Count}(x,l_{1},z_{1})\land\mathit{Count}(x,l_{2},z_{2})\Rightarrow z_{1}=z_{2}$. Thus we have proved the satisfiability of $S_{\mathit{sort}}$. ∎

### v2

@@ 5.1 Constrained Horn Clauses over Words
In Example 5.2 above, if we replace $\mathtt{CTPinf}$ used inside the algorithm of Theorem 3.7 with $\mathtt{TPinf}$, then the inference may fail. Consider the state $M_{2}=\{(\epsilon,ab,ab),(a,b,ab)\}$ above. If $\mathtt{TPinf}$ is invoked instead of $\mathtt{CTPinf}$, we may obtain $t_{2}=(x,y,xy)$ as a minimal (but not the least) $t$ such that $M_{2}\subseteq\mathcal{L}(t)$. Then, as a model of $(l^{\prime}_{1},xl_{2},l_{3})\in\mathcal{L}(t_{2})\land|x|=1\land(xl^{\prime}_{1},l_{2},l_{3})\not\in\mathcal{L}(t_{2})$, we may obtain $(x,l^{\prime}_{1},l_{2},l_{3})=(c,ab,\epsilon,abc)$, which yields $\widetilde{s}_{2}^{\prime}=(cab,\epsilon,abc)$.
By setting $M_{3}^{\prime}=\{(\epsilon,a,a),(a,b,ab),(cab,\epsilon,abc)\}$, we obtain $\mathtt{TPinf}(M_{3}^{\prime})=(x,y,zy)$. Note that $\mathcal{L}(x,y,x^{R}y)\subsetneq\mathcal{L}(x,y,zy)$. Thus, we have failed to infer the least model $\{\mathit{Reva}\mapsto(x,y,x^{R}y)\}$ of $\mathcal{S}_{D}$, and therefore, we have also failed to prove that $\mathcal{S}_{\mathit{Reva}}$ is satisfiable.
We can still prove the satisfiability of $\mathcal{S}_{\mathit{Reva}}$ by using only STPs, based on Theorem 5.2 (which relies on Theorem 3.8 instead of Theorem 3.7). ∎
@@ 5.2 Combination with CHC Solving for Integer Arithmetic
The CHCs obtained from list-manipulating programs typically contain integer constraints in addition to word (or list) constraints. The CSTPs alone are not always sufficient to express appropriate invariants for such CHCs. Let us recall the second example (involving take and drop) in Section 1. It can be expressed as the satisfiability of the following CHCs:
$\mathit{Take}(0,l,\epsilon).\qquad n\neq 0\Rightarrow\mathit{Take}(n,\epsilon,\epsilon).\qquad n\neq 0\land\mathit{Take}(n-1,l^{\prime},r)\land|x|=1\Rightarrow\mathit{Take}(n,xl^{\prime},xr).$
$\mathit{Drop}(0,l,l).\qquad n\neq 0\land\mathit{Drop}(n,\epsilon,\epsilon).\qquad n\neq 0\land\mathit{Drop}(n-1,l^{\prime},r)\Rightarrow\mathit{Drop}(n,xl^{\prime},r).$
@@
Here, we have omitted the clauses for $\mathit{Append}$. As before we have encoded lists into words, but the resulting CHCs also contain integer constraints. Ignoring the integer constraints, the least CSTP model of the definite clauses obtained by $\mathtt{CTPinf}$ is: $\mathit{Take}(n,l_{1},l_{2})\equiv\exists l_{3}.l_{1}=l_{2}l_{3}$, $\mathit{Drop}(n,l_{1},l_{2})\equiv\exists l_{3}.l_{1}=l_{3}l_{2}$, and $\mathit{Append}(l_{1},l_{2},l_{3})\equiv l_{3}=l_{1}l_{2}$. It is not strong enough to satisfy the goal clause (i.e., the last clause above).
To address the issue, we use a CHC solver over integers to strengthen the model of definite clauses. We abstract lists to their lengths and obtain the following abstract version of the CHCs.
@@
In general, given a system of CHCs over lists $\mathcal{S}=\mathcal{S}_{D}\cup\mathcal{S}_{G}$, (i) the list constructors [][\,] and $\mathbin{::}$ are replaced with their length abstractions 0 and $\lambda(x,y)=1+y$, (ii) the list equalities (which implicitly occur in the clause bodies) are replaced with integer equalities; and (iii) the list inequalities $l\neq l^{\prime}$ in the body of goal clauses $\mathcal{S}_{G}$ are replaced with the integer inequalities, and those in the bodies of definite clauses $\mathcal{S}_{D}$ are replaced with $\mathit{true}$.
The above abstraction ensures that a model of the abstract version $\mathcal{S}^{\prime}_{D}$ of definite clauses yields that of the original definite clauses $\mathcal{S}_{D}$: if $P^{\prime}(\widetilde{x},\widetilde{l}^{\prime})\equiv\varphi$ satisfies $\mathcal{S}^{\prime}_{D}$, then $P(\widetilde{x},\widetilde{l})\equiv[|\widetilde{l}|/\widetilde{l}^{\prime}]\varphi$ satisfies $\mathcal{S}_{D}$ (where we assume $\widetilde{x}$ are integer arguments and $\widetilde{l}$ are list arguments); this follows from the soundness of the catamorphism-based abstraction of CHCs [33].
In contrast, there is no such guarantee for the goal clauses: because of the replacement of $_{\mathit{list}}$ and $\neq_{\mathit{list}}$ with $_{\mathit{int}}$ and $\neq_{\mathit{int}}$, the abstract version of goal clauses is neither stronger nor weaker than the original goal clauses. (Notice that $|l_{1}|=|l_{2}|$ is an over-approximation of $l_{1}=l_{2}$, while $|l_{1}|\neq|l_{2}|$ is an under-approximation of $l_{1}\neq l_{2}$.) The abstract goal clauses, however, tend to help a CHC solver to find an appropriate invariant. In the case above, if $l\neq l^{\prime}$ were replaced with $\mathit{true}$, we would obtain $\mathit{Take^{\prime}}(n,l,r_{1})\land\mathit{Drop^{\prime}}(n,l,r_{2})\land\mathit{Append^{\prime}}(r_{1},r_{2},l^{\prime})\Rightarrow\mathit{false}$, which is so strong that the abstract CHCs would become unsatisfiable.
@@
$\mathit{Take^{\prime}}(n,l,r)\equiv r=l<n\lor l\geq n=r\qquad\mathit{Drop^{\prime}}(n,l,r)\equiv(l<n\land r=0)\lor l\geq n=l-r.$
@@
$\mathit{Take}(n,l,r)\equiv|r|=|l|<n\lor|l|\geq n=|r|\qquad\mathit{Drop}(n,l,r)\equiv(|l|<n\land|r|=0)\lor|l|\geq n=|l|-|r|.$
@@
$\mathit{Take}(n,l,r)\equiv\exists s.l=rs\land(|r|=|l|<n\lor|l|\geq n=|r|)$
$\mathit{Drop}(n,l,r)\equiv\exists s.l=sr\land((|l|<n\land|r|=0)\lor|l|\geq n=|l|-|r|).$
@@
Instead of just using the information obtained from the length abstraction, it is possible to propagate information gathered from STP inference and length abstraction to each other, as in Nelson-Oppen procedure [46] for theory combinations. For example, suppose that $P(l_{1},l_{2},l_{3})$ is found to imply $l_{3}=l_{1}l_{2}$ by STP inference. Then, each occurrence of $P(l_{1},l_{2},l_{3})$ in the body of a clause can be replaced by $P(l_{1},l_{2},l_{3})\land l_{3}=l_{1}l_{2}$, before applying the length abstraction. Then, a clause of the form $P(l_{1},l_{2},l_{3})\land\cdots\Rightarrow P(l_{4},l_{5},l_{6})$ is abstracted to $P^{\prime}(l_{1},l_{2},l_{3})\land l_{3}=l_{1}+l_{2}\land\cdots\Rightarrow P^{\prime}(l_{4},l_{5},l_{6})$. ∎
@@ 5.3 Beyond Sequence and Integer Constraints
The (conjunctive) solvable set/multiset tuple patterns discussed in Section 4.2 are also applicable to CHC solving, as stated in the theorem below. Here, CHCs on multisets refer to those whose constraint formulas are drawn from those on multisets [50]. The theorem follows from the set/multiset versions of Theorems 3.7, 3.8, and 3.10.
Theorem 5.4.
Given a system $\mathcal{S}$ of CHCs over multisets, it is decidable whether $\mathcal{S}$ has a CSMTP-model. Similarly, given a system $\mathcal{S}$ of CHCs on sets, it is decidable whether $\mathcal{S}$ has a CSSTP-model.
Theorem 5.5.
Given a system $\mathcal{S}$ of CHCs over multisets, it is decidable whether $\mathcal{S}$ has a SMTP-model. Similarly, given a system $\mathcal{S}$ of CHCs on sets, it is decidable whether $\mathcal{S}$ has a SSTP-model.
@@
Let us consider the following insertion sort program:
  let rec insert x l =
    match l with [] -> [x] | y::l’ -> if x<y then x::l else y::(insert x l’)
  let rec isort l = match l with [] -> [] | x::l’ -> insert x (isort l’)
  let main l = assert(eq_as_multiset (isort l) l).
The main function takes a list $l$ as input, sorts it, and asserts that the result is equivalent to $l$ as multisets; here, for the sake of simplicity, we assume that eq_as_multiset is a primitive. By abstracting lists to multisets and overapproximating the formula $x<y$ by $\mathit{true}$, we obtain the following CHCs on multisets, whose satisfiability implies the lack of assertion failures.
$|x|=1\Rightarrow\mathit{Insert}(x,\epsilon,x).\qquad l=yl^{\prime}\land|x|=|y|=1\Rightarrow\mathit{Insert}(x,l,xl).$
$l=yl^{\prime}\land|x|=|y|=1\land\mathit{Insert}(x,l^{\prime},r)\Rightarrow\mathit{Insert}(x,l,yr).$
$\mathit{Sort}(\epsilon,\epsilon).\qquad|x|=1\land l=xl^{\prime}\land\mathit{Sort}(l^{\prime},r^{\prime})\land\mathit{Insert}(x,r^{\prime},r)\Rightarrow\mathit{Sort}(l,r).\qquad\mathit{Sort}(l,r)\Rightarrow l=r.$
By running the multiset version of the procedure of Theorem 3.7, we obtain the following least CSMTP-model for definite clauses:
$\mathit{Insert}(x,l,r)\equiv(x,l,r)\in\mathcal{L}(x,l,xl)\qquad\mathit{Sort}(l,r)\equiv(l,r)\in\mathcal{L}(l,l).$
Since it satisfies the goal clause $\mathit{Sort}(l,r)\Rightarrow l=r$, we can conclude that the system of CHCs above is satisfiable, and hence also that the program does not suffer from assertion failures. In the same manner, we can also verify that the merge and quick sort programs preserve the multiset of elements, fully automatically (without any invariant annotations for auxiliary functions such as merge and partition). ∎